A small park is in the shape of a rectangle that measures 525 feet by 468 feet.
a. Find the number of feet of fencing that would be needed to enclose the park. Express your answer to the nearest foot.
b. If the entire park is to be painted with grass seed, find the number of square feet to be seeded. Express your answer to the correct number of significant digits based on the given dimensions. : A small park is in the shape of a rectangle that measures 525 feet by 468 feet.
c. The grass seed to be purchased is packaged in sacks, each of which holds enough seed to cover 25,000 square feet of ground. How many sacks of seed are needed to seed the park?

Answer by Alan3354(69443)

A small park is in the shape of a rectangle that measures 525 feet by 468 feet.
a. Find the number of feet of fencing that would be needed to enclose the park. Express your answer to the nearest foot.
Perimeter = 2*Width + 2*Length
P = 2*525 + 2*468 = 1050+936
P = 1986 feet
-----------------
b. If the entire park is to be painted with grass seed, find the number of square feet to be seeded. Express your answer to the correct number of significant digits based on the given dimensions. : A small park is in the shape of a rectangle that measures 525 feet by 468 feet.
Area = L*W = 525*468
A = 245,700 square feet
-------------------
c. The grass seed to be purchased is packaged in sacks, each of which holds enough seed to cover 25,000 square feet of ground. How many sacks of seed are needed to seed the park?
Sacks = Area/25,000 = 245,700/25,000
Sacks = 9.828 so you have to use 10 sacks